# Trapped fireshell (halo) of photons and pairs around black-hole horizon: source for ultra-high-energy particles

### v1

@@ 5.2 UHE emission luminosity and energy spectrum
Section 4 shows that the KN cross section (4.1) decreases as accelerated electron and proton energies increase. Therefore, we expect high-energy electrons to stream out and carry away the halo energy $E_{\gamma}$ of photons and pairs. Using the rates of radiation energy transfer to an accelerated electron and proton pair (4.26), total bulk electron number $\bar{N}_{e}=BN_{\gamma}$ (3.3) and photon-pair number $N_{\gamma}$ (2.10), we obtain the luminosity of high-energy electrons and protons of the Lorentz factor $\gamma_{e}$,
@@
indicating the effective number of bulk electrons $\bar{N}_{e}$ accelerated inside the volume $(\sigma_{{}_{T}}\delta\ell)$ of one-dimensional tube of photon-electron interactions. We treat the $A$ as a parameter, depending on the $B\ll 1$ and $(\sigma_{{}_{T}}\delta\ell m_{e}^{3})=(8\pi/3)\alpha^{2}(\delta\ell/\lambda_{e})$, which can be much larger than one.
@@
    The luminosity sensitively depends on the horizon temperature $T_{\gamma}^{+}$ (2.9), via the photons and pairs number density $n^{+}_{\gamma}\equiv n_{\gamma}(T_{\gamma}^{+})$ (2.8) and the fraction $(N_{e}/\bar{N}_{e})$ (4.15) of electrons accelerated up to the Lorentz factor $\gamma_{e}$. Figure 5 shows that the fraction $N_{e}/\bar{N}_{e}$ is exponentially suppressed by small temperatures $T^{+}_{\gamma}\lesssim 10m_{e}$, therefore, the UHE emission luminosity (5.3) is completely negligible in this case, unless the non-trivial acceleration number $A\gg 1$ (5.4), which relates to the baryon-loading $B$ and size $\delta\ell$ of the horizon halos (spots) of energetic photons and pairs.
@@
    The fraction $N_{e}/\bar{N}_{e}$ properties (4.16) and Figure 5 imply the UHE emission luminosity (5.3) of particles’ energies $m_{e}^{\rm eff}\gamma_{e}$ follows the spectral behaviour
@@
    as shown in Fig. 6. This is consistent with the Compton-rocket effect (3.1) and (3.2) [5], which states that relativistic electrons gain more energy ($\gamma_{e}>1$) from radiation photons than non-relativistic ones ($\gamma_{e}\approx 1$).
@@
    Similar to the blackbody radiation (5.1), the UHE emission is not a transient phenomenon, but rather a lasting, continuous radiation, although the luminosity increases with the UHE particles’ energy $m_{e}^{\rm eff}\gamma_{e}$. The reason is that the fraction $(N_{e}/\bar{N}_{e})$ becomes tiny, and the emission time scale
@@
For the second point (2), the increment luminosity $L_{e}\propto(\gamma_{e})^{2}$ with $\gamma_{e}$ does not run to infinity $L_{e}\propto(\gamma_{e})^{2}\rightarrow\infty$ in time. The reason is that the large luminosity $L_{e}$ decreases the halo temperature $T_{\gamma}^{+}$ and energy $E_{\gamma}$ faster in time, as required by total energy conservation. As a result, the accelerated electron fraction $N_{e}/\bar{N}_{e}$ (4.17) and the UHE emission luminosity $L_{e}$ (5.3) of large $\gamma_{e}$ decreases fast in time, as will be shown below.
To compare two emission processes from the horizon halo, we calculate the ratio of the blackbody radiation (5.1) and UHE emission (5.3) luminosities given by
@@
There are two possibilities: (i) the blackbody radiation is dominant $L_{e}/L^{0}_{\gamma}\ll 1$ for small temperature $T^{+}_{\gamma}\gtrsim m_{e}$ and fraction $N_{e}/\bar{N}_{e}\ll 1$; (ii) the UHE emission is dominant $L_{e}/L^{0}_{\gamma}\gg 1$ for large temperature $T^{+}_{\gamma}\gg m_{e}$ and the fraction $N_{e}/\bar{N}_{e}$ is significant non-trivial. The former case is well studied, and the blackbody luminosity is small due to the halo’s large opacity $(r_{+}/\xi_{\gamma})\gg 1$. We are primarily interested in the latter case for $\gamma_{e}\gg 1$, when $T^{+}_{\gamma}\gg m_{e}$ and the UHE emission dominantly dissipates the energy $E_{\gamma}$ (2.5) of the horizon halo.
Figure 7: For illustrating the numerical results, we adopt the initial horizon temperature $T_{\gamma}^{+}(0)=100m_{e}\approx 50$ MeV (2.9) and the collision number $A=10^{-3}$ (5.4). We consider the Lorentz factor $\gamma_{e}\sim 10^{9}$ channel, corresponding to UHE electron energies $\sim 10^{3}$ TeV and proton energies $\sim 10^{3}$ PeV. The Compton time $\tau_{e}=(\lambda_{e}/c)=1.28\times 10^{-21}{\rm sec}$, and the final time $10^{26}\tau_{e}$ is about days. The $L_{e}/L_{\gamma}^{0}$ broken line (log-log plot) indicates the transition tuning place from the UEH emission dominant $L_{e}/L^{0}_{\gamma}\gg 1$ to the blackbody ($B$) emission dominant $L_{e}/L^{0}_{\gamma}\ll 1$. For different values of $\gamma_{e}$ and $T_{\gamma}^{+}(0)$, the results will exhibit the same behaviours in time variations. The UHE emission luminosity $L_{e}$ will increase for large initial temperature $T_{\gamma}^{+}(0)$ and acceleration number $A$, and the emission will last longer for smaller $\gamma_{e}$ channels of UHE particles.
5.3 UHE emission luminosity and halo temperature decrease in time
The blackbody emission (5.1) and the UHE emission (5.3) dissipate the halo thermal energy $E_{\gamma}$ (2.5), therefore, the energy conservation yields
@@ 5.3 UHE emission luminosity and halo temperature decrease in time
where the factor $\eta$ in the second line comes from the boundary condition (2.3). In the halo decay equation (5.9), we particularly consider a large $\gamma_{e}\gg 1$ channel, because it dominates the energy dissipation of the horizon halo. Before presenting numerical calculations, we note that in the luminosities $L_{\gamma}^{0}$ (5.1) and $L_{e}$ (5.3), and $dT^{+}_{\gamma}/dt$ (5.9), the gravitational redshift in energy and the gravitational dilation in time are simplified.
For a given $\gamma_{e}\gg 1$ value and an initial value $T^{+}_{\gamma}(0)$ at the time $t=0$, we use the numerical result $N_{e}/\bar{N}_{e}$ (4.15) as a function of $T^{+}_{\gamma}(t)$ to integrate the halo decay equation (5.9) over the time. Furthermore, substituting the numerical solution $T^{+}_{\gamma}(t)/m_{e}$ into $(N_{e}/\bar{N}_{e})$ and $(n_{\gamma}^{+}/m_{e}^{3})$ in Eq. (5.3), we calculate the UHE emission luminosity $L_{e}(t)$ declining as a function of the time $t/\tau_{e}$. Although complex, we calculate the numerical integrations using the program in Mathematics. Figure 7 shows that the numerical results of the horizon temperature $T^{+}_{\gamma}(t)/m_{e}$, accelerated electron fraction $N_{e}(t)/\bar{N}_{e}$, and the UHE emission luminosity $L_{e}(t)$ monotonically decrease in the time $t/\tau_{e}$ varying up to $10^{26}\tau_{e}$, which is about days. The results show two phases. (i) The initial phase: the UHE emission initially dominates over the blackbody radiation $L_{e}/L_{\gamma}^{0}>1$, and the turning point to $L_{e}/L_{\gamma}^{0}<1$ occurs around $t=10^{25}\tau_{e}$ when $T^{+}_{\gamma}\approx 10m_{e}$, and $N_{e}(t)/\bar{N}_{e}$ drops significantly. (ii) The later phase: the horizon halo cooling process attributed to the blackbody radiation $L_{\gamma}^{0}$ (5.1) and the temperature $T^{+}_{\gamma}$ drops at a different slope from that in the initial phase.
In the UHE emission luminosity dominant phase $L_{e}/L^{0}_{\gamma}\gg 1$, we neglect the small time-variation effect of the blackbody luminosity $L_{\gamma}^{0}(t)$ (5.1) caused by the time varying mean-free path $\xi_{\gamma}=(\sigma_{\gamma}n^{+}_{\gamma})^{-1}$ (2.7) and number density $n^{+}_{\gamma}\propto[T^{+}_{\gamma}(t)/m_{e}]^{3}$. The effect should probably be considered in the blackbody luminosity $L_{e}/L^{0}_{\gamma}\ll 1$ dominant phase, when the horizon temperature $T^{+}_{\gamma}<10m_{e}$, electron-positron pair annihilation into two photons occurs and the halo opacity drops. As a consequence, the TOV equilibrium (2.2) is lost and the horizon halo (or spot) shrinks and disappears. These are not the topics of this study and will be subjects of future investigations.
We adopt a simplified model to study the acceleration processes of UHE electrons and protons. These processes occur when a fireball (spot) forms during gravitational processes of accretion, collapse and coalescence. The entire process is complex, and numerical studies are required.
@@ 6 Characteristics relevant for observations and experiments
6.1 Peculiar features of the mechanism accelerating UHE particles
Based on a trapped fireshell halo around the horizon, the Compton-rocket effect, and unstable runaway dynamics, we qualitatively discuss a possible microscopic mechanism for producing UHE particles (electrons and protons) in the opaque, high-temperature fluid of photons and pairs, which can form in high-energy astrophysical processes and environments. Discussions and calculations are in an idealised one-dimensional model. Nevertheless, we highlight some qualitative characteristics of this mechanism:
@@ 6.1 Peculiar features of the mechanism accelerating UHE particles
    The mechanism occurs at local microscopic length and time scales, over which the photon-pair fluid temperature ($T_{\gamma}\gg m_{e}$) has a large gradient, resulting in a strong radiation force for electron and proton acceleration. Therefore, the mechanism does not essentially depend on the global symmetry, macroscopic length and time scales of high-energy astrophysical environments and strong gravitational processes, which create such a fire spot of highly dense and energetic fluid of photons and pairs. Therefore, the trapped fireshell (spherical halo) near the horizon is only an example for simplifying calculations and illustrations. The fire spots can be formed, for instance, in gravitational binary coalescence and accretion cases, and may be in future laser experiments.
@@
    The horizon temperature $T^{+}_{\gamma}>10m_{e}$, as if an effective overcritical classical electric field $E_{\rm eff}\sim{\mathcal{O}}(10)E_{c}$, provides a significant probability of accelerating electrons and protons to UHE up to TeV, PeV, EeV scales and even higher. The total luminosity of UHE particles increases as the acceleration number $A$ increases, namely, more electrons in the fireshell participate in the acceleration process. The UHE particle luminosity (5.3) could be significant for the small temperature $2m_{e}<T^{+}_{\gamma}<10m_{e}$ and the large acceleration number $A$. However, we cannot show the results, because the numerical precision is limited in calculating the exponential factor $e^{-m_{\rm eff}/T_{\gamma}^{+}}$.
@@
    The UHE emission is not a transient burst, but a long-lasting continuous radiation. The luminosity of UHE particles has a peculiar spectral behaviour: an approximate power law $\gamma^{-\nu}_{e}$ for $\gamma_{e}\lesssim 10^{3}$, and being proportional to $\gamma_{e}^{2}$ for $\gamma_{e}\gtrsim 10^{3}$. For the appropriate values of the photon and pair temperature $T^{+}_{\gamma}>10m_{e}$ and acceleration number $A$, the resultant UHE emission luminosity (emission probability) and energy spectra (Lorentz factor $\gamma_{e}$ or $\gamma_{p}$) can be relevant for observations.
UHE electrons and protons interact with fields and particles in the medium around their sources, producing VHE photons and neutrinos. Since these interactions are highly energetically forward scattering, resultant secondary VHE particles carry the characteristics of initial primary UHE particles. Therefore, the characteristics (1)-(3) in terms of only two basic parameters $T^{+}_{\gamma}$ and $A$ can be relevant for verification or falsification by the observations of UHE and VHE particles.

### v2

@@ 6.2 UHE emission luminosity and energy spectrum
Section 4 shows that the KN cross section (4.1) decreases as accelerated electron and proton energies increase. Therefore, we expect high-energy electrons to stream out and carry away the fireshell energy $E_{\gamma}$ of photons and pairs. Using (i) the rates of radiation energy transfer to an accelerated electron and proton pair (5.10), (ii) total bulk electron number $\bar{N}_{e}=BN_{\gamma}$ (2.13) and (iii) photon-pair number $N_{\gamma}$ (2.11), we obtain the luminosity of high-energy electrons (protons) of the Lorentz factor $\gamma_{e}\approx\gamma_{p}$,
@@
indicating the effective number of bulk electrons $\bar{N}_{e}$ accelerated inside the volume $(\sigma_{{}_{T}}\delta\ell)$ of one-dimensional tube of photon-electron interactions. We treat the $A$ as a parameter, depending on the baryon matter contamination $B\ll 1$, the firespot size $\delta\ell$ and $(\sigma_{{}_{T}}\delta\ell m_{e}^{3})=(8\pi/3)\alpha^{2}(\delta\ell/\lambda_{e})$, which can be much larger than one.
@@
    The luminosity sensitively depends on the horizon temperature $T_{\gamma}^{+}$ (2.10), via the photons and pairs number density $n^{+}_{\gamma}\equiv n_{\gamma}(T_{\gamma}^{+})$ (2.9) and the fraction $(N_{e}/\bar{N}_{e})$ (4.15) of electrons accelerated up to the Lorentz factor $\gamma_{e}$. Figure 5 shows that the fraction $N_{e}/\bar{N}_{e}$ is exponentially suppressed by small temperatures $T^{+}_{\gamma}\lesssim 10m_{e}$, therefore, the UHE emission luminosity (6.3) is completely negligible in these cases, unless the non-trivial acceleration number $A\gg 1$ (6.4), which relates to the baryon-loading $B$ and size $\delta\ell$ of the fireshells (spots) of energetic photons and pairs.
@@
    The fraction $N_{e}/\bar{N}_{e}$ properties (4.16), (4.17) and Figure 5 imply the UHE emission luminosity (6.3) of particles’ energies $m_{e}^{\rm eff}\gamma_{e}$ follows the spectral behaviour
@@
    as shown in Fig. 6. This is consistent with the Compton-rocket effect (3.1) and (3.2) [21], showing that relativistic electrons gain more energy ($\gamma_{e}>1$) from radiation photons than non-relativistic ones ($\gamma_{e}\approx 1$).
@@
    Similar to the blackbody radiation (6.1), the UHE emission is not a transient phenomenon, but rather a lasting, continuous radiation. The reason is that the fraction $(N_{e}/\bar{N}_{e})$ becomes tiny, and the emission time scale
@@
In the second point (2), the luminosity increases with the UHE particles’ energy $m_{e}^{\rm eff}\gamma_{e}$, but the increment in luminosity $L_{e}\propto(\gamma_{e})^{2}$ with $\gamma_{e}$ does not run to infinity $L_{e}\propto(\gamma_{e})^{2}\rightarrow\infty$ in time. The reason is due to the back reaction that the large luminosity $L_{e}$ of large $\gamma_{e}$ decreases the fireshell temperature $T_{\gamma}^{+}$ and energy $E_{\gamma}$ fast in time, as required by total energy conservation. As a result, the accelerated electron fraction $N_{e}/\bar{N}_{e}$ (4.18) and the UHE emission luminosity $L_{e}$ (6.3) of large $\gamma_{e}$ decrease fast in time, as will be shown below.
To compare two emission processes from the fireshell, we calculate the ratio of the blackbody radiation (6.1) and UHE emission (6.3) luminosities given by
@@
There are two possibilities:
  1. (i)
    the blackbody radiation is dominant $L_{e}/L^{0}_{\gamma}\ll 1$ for low temperature $T^{+}_{\gamma}\gtrsim m_{e}$ and small fraction $N_{e}/\bar{N}_{e}\ll 1$;
  2. (ii)
    the UHE emission is dominant $L_{e}/L^{0}_{\gamma}\gg 1$ for large temperature $T^{+}_{\gamma}\gg m_{e}$ and the significantly non-trivial fraction $N_{e}/\bar{N}_{e}$.
The former case (i) is well studied, and the blackbody luminosity is small due to the fireshell’s large opacity $(r_{+}/\xi_{\gamma})\gg 1$. We are primarily interested in the latter case (ii) for $\gamma_{e}\gg 1$, when $T^{+}_{\gamma}\gg m_{e}$ and the UHE emission dominantly dissipates the energy $E_{\gamma}$ (2.6) of the fireshell.
Figure 7: For illustrating the numerical results, we adopt the initial horizon temperature $T_{\gamma}^{+}(0)=100m_{e}\approx 50$ MeV (2.10) and the acceleration number $A=10^{-3}$ (6.4). We consider the Lorentz factor $\gamma_{e}\sim 10^{9}$ channel, corresponding to UHE electron energies $\sim 10^{3}$ TeV and proton energies $\sim 10^{3}$ PeV. The Compton time $\tau_{e}=(\lambda_{e}/c)=1.28\times 10^{-21}{\rm sec}$, and the final time $10^{26}\tau_{e}$ is about days. The $L_{e}/L_{\gamma}^{0}$ broken line (log-log plot) indicates the transition tuning place from the UEH emission dominant $L_{e}/L^{0}_{\gamma}\gg 1$ to the blackbody ($B$) emission dominant $L_{e}/L^{0}_{\gamma}\ll 1$. For different values of $\gamma_{e}$ and $T_{\gamma}^{+}(0)$, the results will exhibit the same behaviours in time variations. The UHE emission luminosity $L_{e}$ increases for large initial temperature $T_{\gamma}^{+}(0)$ and acceleration number $A$. The emission duration lasts longer for smaller $\gamma_{e}$ channels of UHE particles.
6.3 Fireshell temperature and UHE emission decrease in time
The blackbody emission (6.1) and the UHE emission (6.3) dissipate the fireshell thermal energy $E_{\gamma}$ (2.6), therefore, the total energy conservation yields
@@ 6.3 Fireshell temperature and UHE emission decrease in time
where the factor $\eta$ in the second line comes from the boundary condition (2.4). In the fireshell decay equation (6.9), we particularly consider a large $\gamma_{e}\gg 1$ channel, because it dominates the energy dissipation of the fireshell. Before presenting numerical calculations, we note that the gravitational redshift factor in energy and the gravitational dilation in time are simplified in the luminosities $L_{\gamma}^{0}$ (6.1) and $L_{e}$ (6.3), and the temperature variation $dT^{+}_{\gamma}/dt$ (6.9).
For a given channel $\gamma_{e}\gg 1$ value and an initial value $T^{+}_{\gamma}(0)$ at the time $t=0$, we use the numerical result $N_{e}/\bar{N}_{e}$ (4.15) as a function of $T^{+}_{\gamma}(t)$ to integrate the fireshell decay equation (6.9) over the time. Furthermore, substituting the numerical solution $T^{+}_{\gamma}(t)/m_{e}$ into $(N_{e}/\bar{N}_{e})$ and $(n_{\gamma}^{+}/m_{e}^{3})$ in Eq. (6.3), we calculate the UHE emission luminosity $L_{e}(t)$ declining as a function of the time $t/\tau_{e}$. We calculate these integrations numerically by using the Mathematica program. Figure 7 shows that the numerical results of the horizon temperature $T^{+}_{\gamma}(t)/m_{e}$, the accelerated electron fraction $N_{e}(t)/\bar{N}_{e}$, and the UHE emission luminosity $L_{e}(t)$ monotonically decrease in time $t/\tau_{e}$ varying up to $10^{26}\tau_{e}$, which is about days. The results show two phases:
  1. (i)
    The initial phase: the UHE emission initially dominates over the blackbody radiation $L_{e}/L_{\gamma}^{0}>1$, and the turning point to $L_{e}/L_{\gamma}^{0}<1$ occurs around $t=10^{25}\tau_{e}$ when the temperature $T^{+}_{\gamma}\approx 10m_{e}$, and the accelerated electron fraction $N_{e}(t)/\bar{N}_{e}$ decreases significantly in time.
  2. (ii)
    The latter phase: the fireshell cooling process attributed to the blackbody radiation $L_{\gamma}^{0}$ (6.1) and the temperature $T^{+}_{\gamma}(t)$ decreases in time at a different slope (power law) from the one in the initial phase.
In the UHE emission luminosity-dominant phase $L_{e}/L^{0}_{\gamma}\gg 1$, we neglect the small time-variation effect of the blackbody luminosity $L_{\gamma}^{0}(t)$ (6.1) caused by the time varying mean-free path $\xi_{\gamma}=(\sigma_{\gamma}n^{+}_{\gamma})^{-1}$ (2.8) and number density $n^{+}_{\gamma}\propto[T^{+}_{\gamma}(t)/m_{e}]^{3}$ of photons and pairs in the fireshell. The effect should probably be considered in the blackbody luminosity $L_{e}/L^{0}_{\gamma}\ll 1$ dominant phase, when the horizon temperature $T^{+}_{\gamma}<10m_{e}$, electron-positron pair annihilation into two photons occurs, and the fireshell opacity decreases. As a consequence, the TOV equilibrium (2.3) is gradually lost and the fireshell (or fire spot) shrinks and disappears. These are not the topics of this study and will be subjects of future investigations.
@@ 7 Characteristics relevant for observations and experiments
7.1 Peculiar features of UHE particles produced in such a scenario
We highlight some qualitative characteristics of UHE particles produced by the scenario discussed in previous sections:
@@ 7.1 Peculiar features of UHE particles produced in such a scenario
    The mechanism occurs at local microscopic length and time scales, over which the photon-pair fluid temperature ($T_{\gamma}\gg m_{e}$) has a large gradient, resulting in a strong radiation force for electron and proton acceleration. Therefore, the mechanism does not essentially depend on the global symmetry, macroscopic length and time scales of high-energy astrophysical environments and strong gravitational processes, which create such a fire spot of highly dense and energetic fluid of photons and pairs. These fire spots contain thermal energy $E_{\gamma}$ (2.6) of the radiation fluid of photons and pairs, which are energetic sources for UHE particle productions. Therefore, the trapped fireshell (halo) near the horizon is only an example to simplify calculations and illustrations. The fire spots can be formed, for instance, in gravitational binary coalescence and accretion cases, in the early Universe, and may be observed in future laser experiments.
@@
    The horizon temperature $T^{+}_{\gamma}>10m_{e}$, as if an effective overcritical classical An electric field $E_{\rm eff}\sim{\mathcal{O}}(10)E_{c}$ provides a significant probability of accelerating electrons and protons to UHE, up to TeV, PeV, EeV, and even higher. The total luminosity of UHE particles increases as the acceleration number $A$ increases, namely, more electrons in the fireshell participate in the acceleration process. The UHE particle luminosity (6.3) could be significant for small temperatures $2m_{e}<T^{+}_{\gamma}<10m_{e}$ and large acceleration numbers $A$. However, we cannot show the results, because the numerical precision is limited in calculating the exponential factor $e^{-m_{\rm eff}/T_{\gamma}^{+}}$.
@@
    The UHE emission is not a transient phenomenon, such as a burst, but a long-lasting continuous radiation. UHE particles have peculiar spectral behaviours in high and low energies: approximate power laws $\propto\gamma^{-\nu}_{e}$ for small Lorentz factor $\gamma_{e}\lesssim 10^{3}$, and $\propto\gamma_{e}^{2}$ for larger $\gamma_{e}$. The luminosity of UHE particles should be suppressed for very large $\gamma_{e}$ due to the back reaction of UHE particle production on the sources of fire spots.
  4. (4)
    Two fire-spot parameters are essential for characterising different sources of producing UHE particles:
    $(T^{+}_{\gamma},~A).$ (7.1)
    The temperatures $T^{+}_{\gamma}$ of fire spots of photons and pairs are crucial for UHE particle spectra ($\gamma_{e}$). The acceleration number $A$ of electrons (protons) participating in acceleration dynamics is relevant to the total luminosity of UHE particles. For the appropriate values $T^{+}_{\gamma}$ and $A$ of the fire spots of astrophysical sources, the resultant UHE emission luminosity (emission probability) and energy spectra (Lorentz factor $\gamma_{e}$ or $\gamma_{p}$) can be relevant for the observations of UHE particles. In general, astrophysical fire-spot sources with large (small) $T^{+}_{\gamma}$ and $A$ values produce large (small) luminosity and energy of UHE particles. While the sources with large $T^{+}_{\gamma}$ and $A$ values are less probable. This may explain why the observed UHE particle luminosity monotonically decreases with energy.
In principle, the primary UHE electrons and protons are energetic enough to produce secondary VHE neutral photons and neutrinos through their decays and/or interactions with low-energy particles in the environments of sources of fire spots. Namely, UHE electrons and protons interact with low-energy fields and particles in the medium around their sources, producing VHE photons and neutrinos. Since these interactions are highly energetically forward scattering, resultant secondary VHE particles carry the characteristics of primary UHE particles. This implies that UHE and VHE particle events should be observationally related to each other and correlate with astrophysical sources.